# TSCMamba: Mamba Meets Multi-View Learning for Time Series Classification

### v1

@@ 3 Methodology
In this section, we describe all the steps of our proposed method step by step. The subsections describe the necessary components and procedures of TSCMamba.
Figure 1: Schematic illustration of the proposed method (TSCMamba). This diagram illustrates the architecture of our approach, featuring tango scanning. Here, DP refers to Depth-wise Pooling (DP) and LP refers to Linear Projection (LP). A switch gate selectively activates the utilization of either ROCKET or MLP-derived features. The MLP module, depicted in the bottom right, comprises two layers with an optional dropout mechanism interspersed for regularization
3.1 Spectrogram Representation
We have chosen Continuous Wavelet Transform (CWT) to represent raw signals in spectrograms. CWT potentially surpasses space-time Fourier transform, fast Fourier transform (FFT), and digital wavelet transform (DWT) by providing superior time-frequency localization and multi-resolution analysis. CWT’s adaptable wavelets enhance feature extraction and noise reduction while better-handling edge effects. This makes CWT particularly suitable for analyzing non-stationary signals. CWT’s continuous and detailed representation may offer a significant advantage over the discrete nature of DWT. This renders CWT highly effective for precise time-frequency analysis.
@@ 3.2 Temporal Feature Extraction
Extracting Local Features with Convolutional Kernels in Unsupervised Fashion. Convolutional kernels usually have limited receptive fields, thus focusing on the extraction of local features. Since an MTS dataset may have local features at multiple temporal scales, it is sensible to capture local features within various widths of receptive fields. To this end, we employ the ROCKET approach (Dempster et al., 2020) to extract local features within various local neighborhoods for each channel in an unsupervised fashion. Here, it is to be noted that we only utilize the time domain to extract the kernel-based features, we do not utilize the class labels of the corresponding features. Therefore, our improvement in performance does not solely rely on the ROCKET method, rather it works as a performance booster in certain datasets.
@@
We apply the ROCKET feature extraction method to each channel of length $L$ to form a feature vector of length $X$. We input our training data as a result, with the input tensor of size $B\times D\times L$, we obtain a tensor $\mathcal{V}_{L}\in{\mathcal{R}}^{B\times D\times X}$ that represents the local features. We utilized the sktime implementation of ROCKET to achieve this (Király et al., 2024; Löning et al., 2019).
@@ 3.3 Fusing Multi-View Representations
After obtaining the spectrogram features using CWT and the temporal features at both local and global levels, we fuse these features to effectively exploit the complementary information in these multi-view representations. Through our empirical study, we observe that for many MTS data, either local features or global features in the temporal domain play a dominant role in discriminating between classes for TSC. This observation motivates us to fuse the spectrogram features with either global or local features in the temporal domain. Denote the temporal features by $\mathcal{V}$, which is either $\mathcal{V}_{G}$ or $\mathcal{V}_{L}$. Then, the fused feature map $\mathcal{V}_{W}$ will be calculated as follows:
@@
where $\lambda\geq 0$ is a learnable parameter that determines the balance between the spectrogram features and the temporal features, $1\leq i\leq B$, $1\leq j\leq D$, $1\leq k\leq X$. We set the $\lambda$ as a learnable parameter while the initial value of $\lambda=1.0$. Therefore, the optimal value of $\lambda$ will be determined during the training process. The initial value of $\lambda=1.0$ ensures a balanced focus initially between temporal and spectral domain features.
@@ 3.4 Inferring with Time-Channel Tango Scanning
With the integrated temporal-spectral contextual representations contained in tensor $\mathcal{U}$, we can now learn salient representations to capture important relationships between features, particularly long-term dependencies. To achieve this, we construct tokens by treating each feature vector in $\mathcal{U}$ along the time and channel dimensions as a separate token. Subsequently, we leverage Mamba, a type of SSM, for modeling the token sequences. Mamba is designed for capturing discriminative contents by selectively scanning the token sequences. This selective scan ability allows the model to focus on the most informative parts of the sequence while ignoring less relevant information. By doing so, Mamba can effectively capture long-term dependencies and identify salient features that are most useful for classification.
@@
Vanilla Mamba Block: Inside a Mamba block, two fully-connected layers in two branches calculate linear projections. The output of the linear mapping in the first branch passes through a 1D causal convolution and SiLU activation ${\mathcal{S}}(\cdot)$ (Elfwing et al., 2018), then a structured SSM. The continuous-time SSM maps an input function or sequence $u(t)$ to output $v(t)$ through a latent state $h(t)$:
@@
For SSMs, diagonal $A$ is often used. Mamba makes $B$, $C$, and $\Delta$ linear time-varying functions dependent on the input. In particular, for a token $u$, $B,C\leftarrow Linear_{N}(u)$, and $\Delta\leftarrow softplus(parameter+Linear_{D}(Linear_{1}(u)))$, where $Linear_{p}(u)$ is a linear projection to a $p$-dimensional space, and $softplus$ activation function. Furthermore, Mamba also has an option to expand the model dimension factor $D$ by a controllable dimension expansion factor $E$. Such coefficient matrices enable context and input selectivity properties (Gu & Dao, 2023) to selectively propagate or forget information along the input token sequence based on the current token. Denote the discretization operation by $\Delta=\tau_{\Delta}(parameter+s_{\Delta})$, where $\tau_{\Delta}$ and $s_{\Delta}$ are both functions of the input. For the special case of univariate sequences, the selectivity property has been mathematically proved (Gu & Dao, 2023), as shown in the following:
@@
Unlike the bi-directional Mamba block in Behrouz & Hashemi (2024); Schiff et al. (2024) that uses two separate SSMs with one for forward direction and the other for backward direction, our tango scanning block uses only a single Mamba block. Importantly, for the inverted sequence in our tango scanning module, the output from the Mamba block is not re-inverted back to the original order before performing element-wise addition. In other words, a tango scanning module only involves sequence inversion once. On the contrary, the bi-directional Mamba block (Behrouz & Hashemi, 2024; Schiff et al., 2024) needs to re-invert the output from the vanilla Mamba block. Empirically, we will demonstrate that our tango scanning can effectively update the hidden state variable while maintaining essentially the same memory footprint as the vanilla Mamba block.
@@
where $\oplus$ denotes element-wise addition of matrices, , and $(s^{(t)})^{T}$ is the transpose of $s^{(t)}$ The resultant fused sequence is of size $D\times 3X$.
@@ 4 Experiments and Result Analysis
Datasets Methods
DTW XGBoost Rocket LSTM LSTNet LSSL TCN Trans Re. In. Pyra Auto. Station. FED. ETS. Flow. DLinear LightTS. TimesNet TSCMamba
Ours
@@
73.6 80.05
In this section, we present the results of our experiments on benchmark datasets for time series classification tasks. We evaluate the performance of our proposed method, TSCMamba, and compare it with several state-of-the-art baseline models. The results demonstrate the effectiveness of TSCMamba in handling complex time series classification tasks.
@@ 4.1 Datasets
We evaluated the performance of our proposed method, TSCMamba, on 10 benchmark datasets for time series classification tasks following TimesNet (Wu et al., 2022a). These datasets are commonly used in the literature and are representative of various domains, including image, audio, and sensor data. We present dataset statistics in S-Table 1. These datasets are sourced from a diverse set of domains and contain a diverse range of classes, channels, and time-sequences leading to a robust benchmark for evaluating classification tasks. Moreover, some datasets contain more data in the Test set than the Train set (EC, HW, HB, JV, SCP1, UG) making the time-series classification a harder task. More domain-related information can be found in Bagnall et al. (2018).
@@ 4.2 Experimental Environment
All experiments were conducted using the PyTorch framework (Paszke et al., 2019) with NVIDIA 4$\times$ V100 GPUs (32GB each). The model was optimized using the ADAM algorithm (Kingma & Ba, 2014) with Cross-Entropy loss following TimesNet (Wu et al., 2022a). Moreover, the baseline results are utilized from TimesNet (Wu et al., 2022a) paper for a fair comparison (Same train-test set across the methods). The batch size, epochs, and initial learning rate varied on the datasets for optimal performance. Moreover, the optimization was performed utilizing a cosine-annealing learning rate scheduler. We measure the prediction performance of our method using accuracy metric where larger values indicate better prediction accuracy.
@@
In this study, we evaluate the performance of our proposed method, TSCMamba, against 19 state-of-the-art baselines, encompassing Transformer-based (Wu et al., 2022a; Vaswani et al., 2017; Kitaev et al., 2020; Zhou et al., 2021; Liu et al., 2021; Wu et al., 2021; Liu et al., 2022; Zhou et al., 2022; Woo et al., 2022; Wu et al., 2022b), CNN-based (Franceschi et al., 2019), RNN-based (Hochreiter & Schmidhuber, 1997; Lai et al., 2018; Gu et al., 2022), MLP-based (Zeng et al., 2023; Zhang et al., 2022), and classical machine learning-based methods (Berndt & Clifford, 1994; Chen & Guestrin, 2016; Dempster et al., 2020). Therefore the comparison among these methods following TimesNet (Wu et al., 2022a) provides strong recent baselines from various aspects of machine learning.
@@ 4.3 Predictive Performance Comparison
The comprehensive results are presented in Table 1. Notably, our approach achieves a substantial improvement of 6.45% over the existing best baseline, TimesNet (Wu et al., 2022a), which is a large margin compared to TimesNet’s own improvement of 0.6% over the previous best baseline, Flowformer (Wu et al., 2022b). This notable performance gain solidifies TSCMamba as a strong contender for the TSC task. We plan to release our code and checkpoints for Table 1 publicly  [^1] https://github.com/Atik-Ahamed/TSCMamba. While in Table 1, we present our best-achieved results, we also demonstrate TSCMamba’s generalizability across 5 runs with mean and error-bar (standard-deviation) in S-Figure 4. From Table 1, it is evident that some methods may perform well in some datasets (TimesNet (Wu et al., 2022a) on JV, SA), however, may lack the performance in a large margin on other datasets (TimesNet (Wu et al., 2022a) on EC, HW). On the contrary, our method keeps a balance across the datasets while showing a large improvement in average.

### v2

@@ 3 Methodology
This section describes our proposed method, TSCMamba, whose overall architecture is schematically illustrated in Fig. 1. The architecture and data flow comprise the following key steps: First, we generate global temporal features, localized temporal features, and joint temporal-spectral features from the MTS data (Subsection 3.1 and 3.2). Second, these diverse views are fused to provide rich contexts for subsequent sequence modeling (Subsection 3.3). A switch gate determines whether to fuse the transform-domain features with local features or global patterns in the temporal domain. Third, both the pre-fusion and post-fusion features are fed into an inference engine consisting of two Mamba blocks. This captures long-term dependencies between these features (Subsection 3.4). Fourth, we introduce a tango scanning scheme for each Mamba block to exploit inversion-invariant features, followed by depth-wise pooling. Fifth, final class decisions are made using an MLP to generate class logits (Subsection 3.5). The following subsections detail each component and procedure of TSCMamba.
Figure 1: Schematic illustration of the proposed model TSCMamba. This diagram illustrates the architecture of our approach, featuring tango scanning. Here, DP refers to Depth-wise Pooling (DP) and LP refers to Linear Projection (LP). A switch gate selectively activates the utilization of either ROCKET or MLP-derived features. The MLP module, depicted in the bottom right, comprises two layers with an optional dropout mechanism interspersed for regularization. Here, $B$ = Batch, $L$ = Length, $D$ = Dimension.
3.1 Spectral Representation
We have chosen Continuous Wavelet Transform (CWT) to represent raw signals in the spectral domain. CWT potentially surpasses space-time Fourier transform, fast Fourier transform (FFT), and digital wavelet transform (DWT) by providing superior time-frequency localization and multi-resolution analysis. CWT’s adaptable wavelets enhance feature extraction and noise reduction while better-handling edge effects. This makes CWT particularly suitable for analyzing non-stationary signals. CWT’s continuous and detailed representation may offer a significant advantage over the discrete nature of DWT. This renders CWT highly effective for precise time-frequency analysis.
@@ 3.2 Temporal Feature Extraction
Extracting Local Features with Convolutional Kernels in Unsupervised Fashion. Convolutional kernels usually have limited receptive fields, thus focusing on the extraction of local features. Since an MTS dataset may have local features at multiple temporal scales, it is sensible to capture local features within various widths of receptive fields. To this end, we employ the ROCKET approach [4] to extract local features within various local neighborhoods for each channel in an unsupervised fashion. Here, it is to be noted that we only utilize the time domain to extract the kernel-based features, we do not utilize the class labels of the corresponding features. Therefore, our improvement in performance does not solely rely on the ROCKET method, rather it works as a performance booster in certain datasets.
@@
We apply the ROCKET feature extraction method to each channel of length $L$ to form a feature vector of length $X$. We input our training data as a result, with the input tensor of size $B\times D\times L$, we obtain a tensor $\mathcal{V}_{L}\in{\mathcal{R}}^{B\times D\times X}$ that represents the local features. We utilized the sktime implementation of ROCKET to achieve this [48, 49].
@@ 3.3 Fusing Multi-View Representations
After obtaining the CWT-based spectral features using CWT and the temporal features at both local and global levels, we fuse these features to effectively exploit the complementary information in these multi-view representations. Through our empirical study, we observe that for many MTS data, either local features or global features in the temporal domain play a dominant role in discriminating between classes for TSC. This observation motivates us to fuse the spectral features with either global or local features in the temporal domain. Denote the temporal features by $\mathcal{V}$, which is either $\mathcal{V}_{G}$ or $\mathcal{V}_{L}$. Then, the fused feature map $\mathcal{V}_{W}$ will be calculated as follows:
@@
where $\lambda\geq 0$ is a learnable parameter that determines the balance between the spectral features and the temporal features, $1\leq i\leq B$, $1\leq j\leq D$, $1\leq k\leq X$. We set the $\lambda$ as a learnable parameter while the initial value of $\lambda=1.0$. Therefore, the optimal value of $\lambda$ will be determined during the training process. The initial value of $\lambda=1.0$ ensures a balanced focus initially between temporal and spectral domain features.
@@ 3.4 Inferring with Time-Channel Tango Scanning
With the integrated temporal-spectral contextual representations contained in tensor $\mathcal{U}$, which is processed by a layer normalization for training stability, we can now learn salient representations to capture important relationships between features, particularly long-term dependencies. To achieve this, we construct tokens by treating each feature vector in $\mathcal{U}$ along the time and channel dimensions as a separate token. Subsequently, we leverage Mamba, a type of SSM, for modeling the token sequences. Mamba is designed for capturing discriminative contents by selectively scanning the token sequences. This selective scanning ability allows the model to focus on the most informative parts of the sequence while ignoring less relevant information. By doing so, Mamba can effectively capture long-term dependencies and identify salient features that are most useful for classification.
@@
Vanilla Mamba Block: Inside a Mamba block, two fully-connected layers in two branches calculate linear projections. The output of the linear mapping in the first branch passes through a 1D causal convolution and SiLU activation ${\mathcal{S}}(\cdot)$ [50], then a structured SSM. The continuous-time SSM maps an input function or sequence $u(t)$ to output $z(t)$ through a latent state $h(t)$:
@@
For SSMs, diagonal $A$ is often used. Mamba makes $B$, $C$, and $\Delta$ linear time-varying functions dependent on the input. In particular, for a token $u$, $B,C\leftarrow Linear_{N}(u)$, and $\Delta\leftarrow softplus(parameter+Linear_{D}(Linear_{1}(u)))$, where $Linear_{p}(u)$ is a linear projection to a $p$-dimensional space, and $softplus$ activation function. Furthermore, Mamba also has an option to expand the model dimension factor $D$ by a controllable dimension expansion factor $E$. Such coefficient matrices enable context and input selectivity properties [19] to selectively propagate or forget information along the input token sequence based on the current token. Denote the discretization operation by $\Delta=\tau_{\Delta}(parameter+s_{\Delta})$, where $\tau_{\Delta}$ and $s_{\Delta}$ are both functions of the input. For the special case of univariate sequences, the selectivity property has been mathematically proved [19], as shown in the following:
@@
Unlike the bi-directional Mamba block in [26, 27] that uses two separate SSMs with one for forward direction and the other for backward direction, our tango scanning block uses only a single Mamba block. Importantly, for the inverted sequence in our tango scanning module, the output from the Mamba block is not re-inverted back to the original order before performing element-wise addition. In other words, a tango scanning module only involves sequence inversion once. On the contrary, the bi-directional Mamba block [27] needs to re-invert the output from the vanilla Mamba block. Empirically, we will demonstrate that our tango scanning can effectively update the hidden state variable while maintaining essentially the same memory footprint as the vanilla Mamba block.
@@
where $\oplus$ denotes element-wise addition of matrices, and $(s^{(t)})^{T}$ is the transpose of $s^{(t)}$. The resultant fused sequence is of size $D\times 3X$.
@@ 3.5 Output Class Representation
3.6 Explanation of Our Scanning Mechanism
In this section, we provide a more comprehensive explanation of our scanning protocol. Although Mamba [19] was developed as a distinct architecture, its outstanding performance across various domains can be interpreted through the lens of the attention mechanism [37]. Building on this theoretical framework and drawing inspiration from Ali et al. [51], we analyze the theoretical foundations of the tango scanning mechanism. Consider a sequence of length $L$ along with its corresponding matrices $\bar{A},\bar{B},C$ from Equation 5. With per-channel time-variant system matrices $\bar{A}_{1},\cdots,\bar{A}_{L}$, $\bar{B}_{1},\cdots,\bar{B}_{L}$, and $C_{1},\cdots,C_{L}$, each channel within the selective state-space layers can be processed independently. For simplicity, let us temporarily assume that the input sequence $u$ consists of a single channel. Given the initial condition $h_{0}=0$, unrolling Eq. 5 yields:
$h_{1}=\bar{B}_{1}u_{1},\quad z_{1}=C_{1}\bar{B}_{1}u_{1},\quad h_{2}=\bar{A}_{2}\bar{B}_{1}u_{1}+\bar{B}_{2}x_{2},\quad z_{2}=C_{2}\bar{A}_{2}\bar{B}_{1}u_{1}+C_{2}\bar{B}_{2}u_{2}$ (16)
$h_{t}=\sum_{j=1}^{t}\big{(}\Pi_{k=j+1}^{t}\bar{A}_{k}\big{)}\bar{B}_{j}u_{j},\quad z_{t}=C_{t}\sum_{j=1}^{t}\big{(}\Pi_{k=j+1}^{t}\bar{A}_{k}\big{)}\bar{B}_{j}u_{j}$ (17)
$z=\tilde{\alpha}u,\quad\begin{bmatrix}z_{1}\\ z_{2}\\ \vdots\\ z_{L}\\ \end{bmatrix}=\begin{bmatrix}C_{1}\bar{B}_{1}&0&\cdots&0\\ C_{2}\bar{A}_{2}\bar{B}_{1}&C_{2}\bar{B}_{2}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ C_{L}\Pi_{k=2}^{L}\bar{A}_{k}\bar{B}_{1}\quad&C_{L}\Pi_{k=3}^{L}\bar{A}_{k}\bar{B}_{2}\quad&\cdots\quad&C_{L}\bar{B}_{L}\end{bmatrix}\begin{bmatrix}u_{1}\\ u_{2}\\ \vdots\\ u_{L}\\ \end{bmatrix}$ (18)
The matrix $\tilde{\alpha}\in\mathbb{R}^{L\times L}$ is a function of the input and other relevant parameters. The element at row $i$ and column $j$ captures how $u_{j}$ influences $z_{i}$, and is computed according to:
$\tilde{\alpha}_{i,j}=C_{i}\Big{(}\Pi_{k=j+1}^{i}\bar{A}_{k}\Big{)}\bar{B}_{j}$ (19)
Equations 18 and 19 establish the connection between $\tilde{\alpha}$ and the standard attention matrix, revealing that S6 operates as a variant of causal self-attention. This form of causality is beneficial for next-token prediction tasks, such as in the natural language processing domain for text generation or in time series forecasting tasks, where the current token must not have access to future tokens. However, in our time series classification tasks, this causality constraint is unnecessary since the entire sequence is available during processing.
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{L}$
Figure 2: Regular forward scanning process. The arrow from node $a$ to node $b$ means $a$ depends on $b$.
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{L}$
Figure 3: Flipped sequence scanning process.
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{L}$
Figure 4: Conceptual illustration of the tango scanning mechanism for enhancing inversion invariance in TSC.
Figure 5: Attention matrix visualization for different scanning protocols, illustrating the attended regions. Tango scanning achieves the maximum possible pairwise attention coverage with one Mamba.
Let $u$ be a time series of length $L$. When using only the regular Mamba, a causal attention mask [51] is applied (Figure 2), restricting the current token $u_{i}$ from attending to any future token $u_{i+1}$. To capture bidirectional dependencies, we introduce a flip-based scanning approach (Figure 3), where each token $u_{i}$ attends to future tokens $u_{i+j}$ with $j\geq 1$. By combining both scanning directions with shared parameters (Figure 4), we achieve full pairwise attention using a single Mamba block without introducing additional parameters. Although this approach may introduce additional calculations for self-terms, it ensures maximum possible attention coverage, leaving no token unattended. We illustrate this strategy for a sequence length using token tango scanning, which can also be extended to channel tango scanning, where each channel is treated as a sequence length and each length as a channel. The practical attention matrix visualization in Figure 5 demonstrates that tango scanning successfully captures all possible token pair interactions.
@@ 4 Experiments and Result Analysis
Dataset Methods
TSCMamba DTW XGBoost Rocket LSTM LSTNet LSSL TCN Trans Re. In. Pyra Auto. Station. FED. ETS. Flow. DLinear LightTS. TimesNet TSLANet
[1994]
[2016]
[2020]
[1997]
[2018]
[2022]
[2019]
[2017]
[2020]
[2021]
[2021]
[2021]
[2022]
[2022]
[2022]
[2022]
[2023]
[2022]
[2023]
[2024]
@@
73.6
76.04
Rank
4.35
15.20
15.55
10.25
19.55
10.40
11.70
12.40
9.40
9.95
8.90
12.80
11.50
7.30
12.40
12.85
6.45
14.60
12.65
6.40 6.40
In this section, we present the results of our experiments on benchmark datasets for time series classification tasks. We evaluate the performance of our proposed method, TSCMamba, and compare it with SOTA baseline models. Our method demonstrates superior performance across datasets of varying complexities. For example, TSCMamba outperforms many baseline methods on datasets with irregular patterns (e.g., EC and HW) in a large margin, indicating its ability to model complex patterns effectively. Furthermore, our approach maintains strong performance even on datasets where traditional models, such as DTW and XGBoost, have historically performed well, demonstrating its versatility. Through comprehensive comparisons with SOTA baselines, we validate the effectiveness of TSCMamba in handling diverse time series classification challenges.
@@ 4.1 Datasets
We evaluated the performance of our proposed method, TSCMamba, on 10 benchmark datasets for time series classification tasks following TimesNet [45] and an additional 20 datasets following TSLANet [8] and UEA archive [53]. These datasets are commonly used in the literature and are representative of various domains, including image, audio, and sensor data. We present dataset statistics in S-Table 1 with more details in Appendix Subsection 9.6. By evaluating our method on all 30 UEA datasets, we ensure a comprehensive assessment across diverse real-world time series classification challenges. This expansion strengthens the generalizability of our findings and aligns with prior SOTA benchmarks. These datasets are sourced from a diverse set of domains and contain a diverse range of classes, channels, and time-sequences leading to a robust benchmark for evaluating classification tasks. Moreover, some datasets contain more data in the Test set than the Train set (EC, HW, HB, JV, SCP1, UG) making the time-series classification a harder task. More domain-related information can be found in [53].
@@ 4.2 Experimental Environment
All experiments were conducted using the PyTorch framework [54] with NVIDIA 4$\times$ V100 GPUs (32GB each). The model was optimized using the ADAM algorithm [55] with Cross-Entropy loss following TimesNet [45]. Moreover, the baseline results are utilized from TimesNet [45] paper for a fair comparison (Same train-test set across the methods). The batch size, epochs, and initial learning rate varied on the datasets for optimal performance. The hyperparameters were selected based on the Train and Test set provided by the dataset archive following TimesNet [45]. Moreover, the optimization was performed utilizing a cosine-annealing learning rate scheduler. We measure the prediction performance of our method using accuracy metric where larger values indicate better prediction accuracy.
@@
In this study, we evaluate the performance of our proposed method, TSCMamba, against 20 state-of-the-art baselines in Table 1, encompassing Transformer-based [8, 45, 37, 38, 52, 42, 41, 39, 44, 40, 43], CNN-based [5], RNN-based [35, 36, 47], MLP-based [30, 31], and classical machine learning-based methods [28, 29, 4]. Therefore the comparison among these methods following TimesNet [45] provides strong recent baselines from various aspects of machine learning.
@@ 4.3 Predictive Performance Comparison
The comprehensive results are presented in Table 1. Notably, our approach achieves a substantial improvement of 4.01% over the existing best baseline, TSLANet [8]. Additionally, it improves upon the existing second-best baseline, TimesNet [45], by 6.45%, which is a large margin compared to TimesNet’s improvement of 0.6% over the previous best baseline, Flowformer [43]. This notable performance gain establishes TSCMamba as a strong contender for the TSC task. Unlike prior incremental improvements in TSC models, the large margin of improvement highlights the effectiveness of our multi-view fusion and tango scanning strategies, reinforcing their impact on long-range dependency modeling. For more insights regarding our scanning mechanism refer to 3.6. Our implementation is available at this link.
While Table 1 presents our best-achieved results, we also demonstrate TSCMamba’s reproducibility and stability across 5 runs with mean and error bars (standard deviation) in S-Figure 11. In addition to the main baselines, we also compare TSCMamba against regular Mamba (Figure 6). TSCMamba consistently outperforms regular Mamba, showcasing the effectiveness of our modifications. Specifically, the integration of tango scanning enhances inter-token dependencies, while multi-view fusion allows the model to adaptively leverage both spectral and temporal features. Beyond outperforming regular Mamba, TSCMamba also demonstrates superior performance compared to CNN-based methods such as TCN, which struggle to maintain sequence-level contextual information. This is evident in datasets where temporal dependencies play a crucial role, further validating the necessity of long-sequence modeling capabilities in TSC tasks. From Table 1, it is evident that some methods may perform well on certain datasets (e.g., TimesNet [45] on JV, SA), but may lack performance by a large margin on others (e.g., TimesNet [45] on EC, HW). In contrast, our method maintains a balance across the datasets while showing a significant improvement in average performance.
In Table 1, we compared our approach with various baselines including traditional methods, where TSLANet [8] achieved the second-best performance. To further validate our model and findings, we evaluated an additional 20 datasets in Table 2 using the recent baselines from TSLANet. Following TimesNet [45] and other recent baselines like TSLANet [8], we present the results as average performance across datasets. However, recognizing that averaging over various datasets might not be the optimal way to assess different models, we also present ranks of different models as suggested by Demšar [56]. These ranks are calculated following the methodology of Ismail Fawaz et al. [6]. This ranking method ensures a more robust comparison by normalizing performance across datasets of varying difficulty levels, providing a clearer picture of relative model performance than averaged accuracy values alone. For both sets of benchmark datasets, Table 1 and Table 2 demonstrate that TSCMamba achieves the best performance in terms of averaged accuracy and rank. Additionally, we provide Critical Difference (CD) diagrams in Appendix Subsection 9.5 to visually compare model rankings across datasets. The diagrams illustrate the relative positioning of TSCMamba alongside other baselines, reinforcing its strong ranking across both benchmark and additional datasets. These visualizations complement our tabular results, providing an alternative perspective on model performance.
Table 2: Additional classification results on the UEA datasets in terms of accuracy (as %). The ranks are computed by averaging dataset-wise rankings of classifiers (lower is better).
Dataset TSCMamba TSLANet GPT4TS TimesNet ROCKET CrossF. PatchTST MLP TS-TCC TS2VEC
(Ours) [2024] [2023] [2023] [2020] [2023] [2023] [2023] [2021] [2022]
AtrialFibrillation 67.00 40.00 33.33 33.33 20.00 46.66 53.33 46.66 33.33 53.33
BasicMotions 100.00 100.00 92.50 100.00 100.00 90.00 92.50 85.00 100.00 92.50
Cricket 98.61 98.61 8.33 87.50 98.61 84.72 84.72 91.67 93.06 65.28
FingerMovements 69.00 61.00 57.00 59.38 61.00 64.00 62.00 64.00 44.00 51.00
HandMovementDirection 71.62 52.70 18.92 50.00 50.00 58.11 58.11 58.11 64.86 32.43
MotorImagery 62.00 62.00 50.00 51.04 53.00 61.00 61.00 61.00 47.00 47.00
PenDigits 98.54 98.94 97.74 98.19 97.34 93.65 99.23 92.94 98.51 97.40
PhonemeSpectra 24.66 17.75 3.01 18.24 17.60 7.55 11.69 7.10 25.92 8.23
RacketSports 91.45 90.79 76.97 82.64 86.18 81.58 84.21 78.95 84.87 74.34
StandWalkJump 73.33 46.67 33.33 53.33 46.67 53.33 60.00 60.00 40.00 46.67
InsectWingbeat 14.30 10.00 10.00 10.00 10.00 10.00 10.00 10.00 10.00 10.00
DuckDuckGeese 62.00 24.00 50.00 56.00 52.00 44.00 24.00 62.00 38.00 68.00
NATOPS 94.44 95.56 91.67 81.82 83.33 88.33 96.67 93.89 96.11 82.22
Libras 90.00 92.78 79.44 77.84 83.89 76.11 81.11 73.33 86.67 85.56
ArticularyWordRecognition 97.00 99.00 93.33 96.18 99.33 98.00 97.67 97.33 98.00 87.33
Epilepsy 97.10 98.55 85.51 78.13 98.55 73.19 65.94 60.14 97.10 62.32
LSST 60.30 66.34 46.39 59.21 54.10 42.82 67.80 35.77 49.23 39.01
ERing 91.50 95.56 95.93 94.07 94.07 94.07 92.59 91.48 90.40 87.40
CharacterTrajectories 99.00 98.75 98.26 97.98 98.75 98.33 97.35 97.35 98.50 99.50
EigenWorms 87.00 41.22 48.09 58.78 78.63 54.96 54.20 41.22 77.90 84.70
Average 77.44 69.51 58.49 67.18 69.15 66.02 67.71 65.40 68.67 63.71
Rank 2.48 3.95 7.40 6.00 4.95 6.17 5.12 6.80 5.22 6.90